OpenStudy (babynini):

Help! Unit Circles. The point p is on the unit circle. Find p(x, y) from the given information. The x coordinate of p is -sqroot6/4 and p lies below the x-axis. p(x,y) =

OpenStudy (anonymous):

Well, you can use circle formula.

OpenStudy (anonymous):

The equation for a circle is: \[ (x-h)^2+(y-k)^2=r^2 \]In this case, \((h,k)\) the center is at \((0,0)\) and \(r\) the radius is \(1\).\[ x^2+y^2=1 \]Now we know that the point lies below the \(x\) axis, which means \(y<0\). And we also know that \(x = -\sqrt{6}/4\)

OpenStudy (anonymous):

Solving for \(y\) gives us: \[ y = \sqrt{1-x^2} = \sqrt{1-\left(\sqrt{6/4}\right)^2} \]

OpenStudy (babynini):

So what would the final answer look like? p= \[(-\sqrt{6}/4, y) \] what is y exactly??

OpenStudy (babynini):

@wio

OpenStudy (babynini):

Oh, or is it p times (x,y) because it says look for p(x,y) ._.

jimthompson5910 (jim_thompson5910):

Is the x coordinate of point p the expression \(\LARGE \frac{-\sqrt{6}}{4}\) ??? OR Is the x coordinate of point p the expression \(\LARGE -\sqrt{\frac{6}{4}}\) ???

OpenStudy (babynini):

the first one

jimthompson5910 (jim_thompson5910):

So if \[\Large x = \Large \frac{-\sqrt{6}}{4}\] then what is x^2 equal to?

OpenStudy (babynini):

\[(-\sqrt{6}/64^2\]?

OpenStudy (babynini):

ah! sorry that got messed up.

OpenStudy (babynini):

\[(-\sqrt{6}/4)^2\]

jimthompson5910 (jim_thompson5910):

good, can you simplify that further?

OpenStudy (babynini):

\[\sqrt{36}/16\] would that be correct?

OpenStudy (babynini):

or..is the 36 still negative?

jimthompson5910 (jim_thompson5910):

If you square \[\Large \sqrt{6}\], the square root will go away. This is because the square and square root undo each other

jimthompson5910 (jim_thompson5910):

or you can think of it like \[\Large \sqrt{6*6}=\sqrt{36} = 6\]

jimthompson5910 (jim_thompson5910):

either way, you'll have 6 in the numerator

OpenStudy (babynini):

Oh gosh. I knew that *facepalm. So then what do I do with the denominator?

jimthompson5910 (jim_thompson5910):

you'll have 16 in the denominator since 4*4 = 16. So that is correct

jimthompson5910 (jim_thompson5910):

x^2 = 6/16 = 3/8 make sure to reduce as much as possible

jimthompson5910 (jim_thompson5910):

now you'll use this to find the value of y the unit circle is the equation x^2 + y^2 = 1 replace x^2 with 3/8 to get (3/8) + y^2 = 1 now you need to solve for y

OpenStudy (babynini):

by putting y under a sq root?

jimthompson5910 (jim_thompson5910):

how would you move that 3/8 over?

OpenStudy (babynini):

\[\sqrt{y}=1-(3/8)\]

OpenStudy (babynini):

or times both sides by 8?

jimthompson5910 (jim_thompson5910):

You did it right. You subtract 3/8 from both sides.

jimthompson5910 (jim_thompson5910):

1-(3/8) = ???

OpenStudy (babynini):

5/8!

jimthompson5910 (jim_thompson5910):

so we know that \[\Large y^2 = \frac{5}{8}\]

jimthompson5910 (jim_thompson5910):

the ultimate goal is to find y itself

OpenStudy (babynini):

so you square root it.

OpenStudy (babynini):

?

OpenStudy (babynini):

\[\sqrt{y^2}=\sqrt{5}/8\]

jimthompson5910 (jim_thompson5910):

the square root will apply to the entire number 5/8 and not just the 5

jimthompson5910 (jim_thompson5910):

so \[\Large y = \pm \sqrt{\frac{5}{8}}\] we are told that "p lies below the x axis" so y must be negative, which means there is a negative sign outside the square root \[\Large y = -\sqrt{\frac{5}{8}}\] notice how the plus/minus turned to just a minus

OpenStudy (babynini):

Okay. So it's p\[(-\sqrt{6}/4,-\sqrt{5/8})\]

OpenStudy (babynini):

?

jimthompson5910 (jim_thompson5910):

Correct, the point is \[\Large \left(\frac{-\sqrt{6}}{4}, -\sqrt{\frac{5}{8}}\right)\]

OpenStudy (babynini):

So when it asks for p(x,y) I'm just meant to find the point and not multiply the p in..right? That would be a whole other thing xD

jimthompson5910 (jim_thompson5910):

p(x,y) is basically the same as saying "p is the point in the form (x,y)" it's unfortunate that it looks like "p times (x,y)" when that's not what it means

OpenStudy (babynini):

Ah, okay. Sir Thompson, you have been absolutely amazing. Thank you so so much! you're a life saver.

OpenStudy (babynini):

Sorry i'm so slow xD tested into a math class so i'm a bit confused as to what all is going on.

jimthompson5910 (jim_thompson5910):

you did fine, so no worries

OpenStudy (babynini):

Thanks again ^-^

jimthompson5910 (jim_thompson5910):

you're welcome